# Error Bounds For Gaussian Process Regression Under Bounded Support Noise With Applications To Safety Certification

### v1

@@ 1 Introduction
Figure 1: Predictive mean and error bounds when learning from 20 samples of $y=x\sin(x)+v$ with $|v|<0.5$. In (a) we plot the comparison of deterministic error bounds and mean predictions, and in (b) we show the probabilistic bounds that hold with 95% probability. We set $\sigma_{n}=0.1$ for our predictions, $\sigma_{n}=0.5$ for Hashimoto et. al. as per Lemma 2, and $\sigma_{n}^{2}=1+2/20$ for Chowdhury et. al. as per Lemma 1. Note bounds for (Abbasi-Yadkori 2013) are nearly identical to Chowdhury et. al. in this example, see Appendix.
@@
Works by Srinivas et al.; Abbasi-Yadkori; Chowdhury and Gopalan; Hashimoto et al. provide formal regression error bounds for GP models under non-Gaussian noise. Specifically, work (Srinivas et al. 2012) first develops a framework for GP regression under the assumption that noise is $R$-sub-Gaussian and then formally quantifies probabilistic bounds on the GP prediction errors. Using a similar framework, (Chowdhury and Gopalan 2017) derives tighter probabilistic bounds. Work (Abbasi-Yadkori 2013) derives similar bounds for Kernel Ridge Regression, which uses the same mean prediction as a GPR without a posterior variance. While applicable, the setting in each of these works is more general than ours, which results in larger error terms when assessed with bounded support noise. Work (Hashimoto et al. 2022) specifically focuses on bounded noise and derives a deterministic error bound. This bound is generally tighter than the probabilistic ones in (Abbasi-Yadkori 2013; Chowdhury and Gopalan 2017); nevertheless, it becomes loose as the size of the support increases. Since both error bounds in (Chowdhury and Gopalan 2017; Hashimoto et al. 2022) are directly applicable to our scenario, we show their derivation in Section 2 and compare our results against them. In the experiments, we also compare against (Srinivas et al. 2012; Abbasi-Yadkori 2013).
@@ 2 Setup and Background
Lemma 1 provides probabilistic bounds on the GP regression error $|\mu_{D}(x)-f(x)|$ by relying on the kernel parameters such as RKHS norm bound $B$ and information gain $\Gamma$, which are often difficult to obtain accurately. They can however be (conservatively) bounded using techniques introduced in  (Srinivas et al. 2012; Jackson et al. 2021a; Hashimoto et al. 2022). Another important observation in Lemma 1 is that, when noise $\mathbf{v}$ has a $R$-sub-Gaussian distribution, variance $\sigma_{n}^{2}$ is set to $1+2/m$, which adds conservatism to the bounds by substantially increasing the variance of GP predictions. Note that to obtain estimates with confidence $1$, i.e., $\delta=0,$ Lemma 1 requires infinitely-many samples. Alternative probabilistic bounds in this setting are considered by Srinivas et al., which makes use of similar parameters, and by Abbasi-Yadkori which leaves $\sigma_{n}$ as a decision variable (see the Appendix). As shown in our experiments, they are also conservative.
@@ 3 Bounded Support Error Bounds
In this section, we introduce novel probabilistic and deterministic error bounds for GP regression for System (1), where noise has a bounded support distribution. We show that in this setting, the results of Lemmas 1 and 2 can be substantially improved.
@@
The proof of Lemma 3 is provided in the Appendix. In Theorem 1, we rely on Hoeffding’s Inequality (Mohri, Rostamizadeh, and Talwalkar 2018) to bound the second term in Eqn (7), which in turn provides a probabilistic bound on the GP regression error when combined with Lemma 3.
@@
The proof uses Lemma 3 for the first term of $\epsilon(x,\delta)$ and derives the second term by applying Hoeffding’s Inequality to $|W_{x}\mathcal{V}|$ by noting that $\mathbb{E}[W_{x}\mathcal{V}]=0$ and each $-\sigma_{v}\leq v_{i}\leq+\sigma_{v}$, enabling the random variable to maintain bounded support on each term. A detailed proof can be found in the Appendix. Note that Theorem 1 only requires two parameters in its probabilistic bound: $c^{*}$ and $B$. In fact, $c^{*}$ can be found by solving the following quadratic optimization problem, where we rely on the boundedness of the support of the distribution of $\mathcal{V}=[v_{1},\ldots,v_{m}]^{T}$:
@@
The other parameter, $B$, can be formally bounded by the technique introduced in (Jackson et al. 2021a). This is in contrast with Lemma 1, which also requires an approximation on the information gain of the kernel.
We should also emphasize that Theorem 1 allows $\sigma_{n}$ to remain as a decision variable, enabling an optimization over $\sigma_{n}$ that can further minimize the error bounds as compared to Lemmas 1 and 2. Recall $G=(K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{2}I)^{-1}$ and that adding more data to $K_{\mathcal{X},\mathcal{X}}$ tends to increase complexity and hence the magnitude of the eigenvalues. This implies that the eigenvalues of $G$ tend to decrease with more data, hence $G^{2}$ then decreases even more rapidly and $K_{x,\mathcal{X}}G^{2}K_{\mathcal{X},x}$ approaches 0 as more data is added. Intuitively, as $m\rightarrow\infty$ then $\lambda_{x}\rightarrow 0$ which implies that we can set $\delta$ arbitrarily close to 0, reducing the bound to the result of Lemma 3 which decreases with samples as $\sigma_{D}(x)$ decreases and $c^{*}$ increases. This further demonstrates an $O(\sqrt{m})$ improvement over the results of Lemma 2.
@@ 4 Application to Safety of Control Systems
The bounds we derive in Theorems 1 and 2 can be particularly important to provide safety guarantees for dynamical systems learnt from data. Specifically, consider the following model of a discrete-time stochastic system with additive noise
@@
A key challenge in many applications is to guarantee that System (13) will stay within a given safe set $X_{s}\subset\mathbb{R}^{d}$, i.e., it avoids obstacles, for a given time horizon $[0,N]$. A common technique to obtain such guarantees is to rely on stochastic barrier functions (Kushner 1967; Santoyo, Dutreix, and Coogan 2021), which represents a generalization of Lyapunov functions to provide set invariance. The intuition behind barrier functions is to build a function $\mathcal{B}:\mathbb{R}^{d}\to\mathbb{R}$ that when composed with System (13) forms a supermartingale and consequently, martingale inequalities can be employed to study the time evolution of the system, as illustrated in the following Proposition.
@@
where $\mu_{D}^{(i)}$ denotes the mean prediction for $f^{(i)}$ and $\overline{\epsilon}_{q}^{(i)}(\delta)$ can be computed by combining Corollary 1 with either Theorem 1 or 2. In the former case, the safety probability in Proposition 1 holds with the confidence resulting from Theorem 1, while in the latter case results hold with confidence $1$. In Section 5, we consider both cases.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our bounds in comparison to the existing state-of-the-art bounds. We first provide a visual of the trends of our bounds as more data is utilized for posterior predictions , as well as demonstrating the improvements that DKL can provide over standard GPs. We then demonstrate how our bounds perform in comparison to the state-of-the-art across several dynamical systems of interest. Finally, we show how our bounds can be used in safety certification, through the use of stochastic barrier functions, as described in Section 4. We note that deep kernels in general are only positive semi-definite, which prohibits the use of techniques in (Maddalena, Scharnhorst, and Jones 2021) for generating a deterministic bound. Further detail on the models considered can be seen in the Appendices.
@@
Safety Certification of Unknown Stochastic Systems We consider the Inverted Pendulum (2D) agent from the Gymnasium environment and a contractive linear 4D system for data-driven safety certification using the formulation in Section 4. We collected data for the Pendulum model under the best controller available from the OpenAI Leaderboard (OpenAI 2024) and perturb the systems with PERT distributions (a transformation of a Beta distribution with a specified mean). We construct a barrier using the Dual-Approach suggested in (Mazouz et al. 2024) and compare results when generating interval bounds on $f$ with Corollary 1 using Theorem 2, Lemma 2, and Theorem 1 with $\delta=0.05$, and using DKL models.
@@ 6 Conclusion
In this paper, we derive novel error bounds (both probabilistic and deterministic) for GP regression under bounded support noise. We demonstrate that by assuming sample noise has zero mean, error bounds that are tighter than sample noise can be achieved. We show that our error bounds utilize significantly more information about the kernel than existing probabilistic bounds; hence, they are well-suited for modern regression techniques that are based on informed kernels such as deep kernel learning. We also show the improvements that our bounds can provide over the state-of-the-art in safety certification of control systems through the use of stochastic barrier functions, generating certificates with significantly larger safety probabilities. Future directions include application of these bounds in safe control and shield design in reinforcement learning.
Appendix A Appendix

### v2

@@ 1 Introduction
Figure 1: Predictive mean and error bounds when learning from 20 samples of $y=x\sin(x)+v$ with $|v|<0.5$. In (a) we plot the comparison of deterministic error bounds and mean predictions, and in (b) we show the probabilistic bounds that hold with 95% probability. We set $\sigma_{n}=0.1$ for our predictions, $\sigma_{n}=0.5$ for Hashimoto et al. as per Lemma 2, and $\sigma_{n}^{2}=1+2/20$ for Chowdhury et al. as per Lemma 1. Note bounds for Abbasi-Yadkori (2013) are nearly identical to Chowdhury et al. in this example, see Appendix A.
@@
Works by Srinivas et al.; Abbasi-Yadkori; Chowdhury and Gopalan; Hashimoto et al. provide formal regression error bounds for GP models under non-Gaussian noise. Specifically, work (Srinivas et al. 2012) first develops a framework for GP regression under the assumption that noise is $R$-sub-Gaussian and then formally quantifies probabilistic bounds on the GP prediction errors. Using a similar framework, (Chowdhury and Gopalan 2017) derives tighter probabilistic bounds. Work (Abbasi-Yadkori 2013) derives similar bounds for Kernel Ridge Regression, which uses the same mean prediction as a GPR without a posterior variance. While applicable, the setting in each of these works is more general than ours, which results in larger error terms when assessed with bounded support noise. Work (Hashimoto et al. 2022) specifically focuses on bounded noise and derives a deterministic error bound. This bound is generally tighter than the prior probabilistic bounds; nevertheless, it becomes loose as the size of the support increases. Since both error bounds in (Chowdhury and Gopalan 2017; Hashimoto et al. 2022) are directly applicable to our scenario, we show their derivation in Section 2 and compare our results against them. In the experiments, we also compare against (Srinivas et al. 2012; Abbasi-Yadkori 2013).
@@ 2 Setup and Background
Lemma 1 provides probabilistic bounds on the GP regression error $|\mu_{D}(x)-f(x)|$ by relying on the kernel parameters such as RKHS norm bound $B$ and information gain $\Gamma$, which are often difficult to obtain accurately. They can however be (conservatively) bounded using techniques introduced in  (Srinivas et al. 2012; Jackson et al. 2021a; Hashimoto et al. 2022). Another important observation in Lemma 1 is that, when noise $\mathbf{v}$ has a $R$-sub-Gaussian distribution, variance $\sigma_{n}^{2}$ is set to $1+2/m$, which adds conservatism to the bounds by substantially increasing the variance of GP predictions. Note that to obtain estimates with confidence $1$, i.e., $\delta=0,$ Lemma 1 requires infinitely-many samples. Alternative probabilistic bounds in this setting are considered by Srinivas et al., which makes use of similar parameters, and by Abbasi-Yadkori which leaves $\sigma_{n}$ as a decision variable (see Appendix A). As shown in our experiments, they are also conservative.
@@ 3 Bounded Support Error Bounds
In this section, we introduce novel probabilistic and deterministic error bounds for GPR for System (1), where noise has a bounded support distribution. We show that in this setting, the results of Lemmas 1 and 2 can be substantially improved.
@@
The proof of Lemma 3 can be found in Appendix B. In Theorem 1, we rely on Hoeffding’s Inequality (Mohri, Rostamizadeh, and Talwalkar 2018) to bound the second term in Eqn (7), which in turn provides a probabilistic bound on the GP regression error when combined with Lemma 3.
@@
The proof uses Lemma 3 for the first term of $\epsilon(x,\delta)$ and derives the second term by applying Hoeffding’s Inequality to $|W_{x}\mathcal{V}|$ by noting that $\mathbb{E}[W_{x}\mathcal{V}]=0$ and each $-\sigma_{v}\leq v_{i}\leq+\sigma_{v}$, enabling the random variable to maintain bounded support on each term. The full proof can be found in Appendix B. Note that Theorem 1 only requires two parameters in its probabilistic bound: $c^{*}$ and $B$. In fact, $c^{*}$ can be found by solving the following quadratic optimization problem, where we rely on the boundedness of the support of the distribution of $\mathcal{V}=[v_{1},\ldots,v_{m}]^{T}$:
@@
The other parameter, $B$, can be formally bounded by the technique introduced in (Jackson et al. 2021a). This is in contrast with Lemma 1, which also requires an approximation on the information gain of the kernel. We should also emphasize that Theorem 1 allows $\sigma_{n}$ to remain as a decision variable, enabling an optimization over $\sigma_{n}$ that can further minimize the error bounds as compared to Lemmas 1 and 2.
We also note as $m\rightarrow\infty$ then $\lambda_{x}$ tends toward 0 which implies that we can set $\delta$ arbitrarily close to 0, reducing the error bound to the result of Lemma 3 which decreases with samples as $\sigma_{D}(x)$ decreases and $c^{*}$ increases. This demonstrates an $O(\sqrt{m})$ improvement over the results of Lemma 2. A detailed discussion is provided in Appendix B.
@@ 4 Application to Safety of Control Systems
The bounds we derive in Theorems 1 and 2 can be particularly important to provide safety guarantees for dynamical systems learned from data. Specifically, consider the following model of a discrete-time stochastic system with additive noise
@@
A key challenge in many applications is to guarantee that System (13) will stay within a given safe set $X_{s}\subset\mathbb{R}^{d}$, i.e., it avoids obstacles, for a given time horizon $[0,N]$. A common technique to obtain such guarantees is to rely on stochastic barrier functions (Kushner 1967; Santoyo, Dutreix, and Coogan 2021), which represents a generalization of Lyapunov functions to provide set invariance. The intuition behind barriers is to build a function $\mathcal{B}:\mathbb{R}^{d}\to\mathbb{R}$ that when composed with System (13) forms a supermartingale and consequently, martingale inequalities can be used to study the systems’ evolution, as shown in the following Proposition.
@@
where $\mu_{D}^{(i)}$ denotes the mean prediction for $f^{(i)}$ and $\overline{\epsilon}_{q}^{(i)}(\delta)$ can be computed by combining Corollary 1 with either Theorem 1 or 2. Work (Mazouz et al. 2024) shows how to calculate a barrier with such constraints. In the former case, the safety probability in Proposition 1 holds with the confidence resulting from Theorem 1, while in the latter case results hold with confidence $1$. In Section 5, we consider both cases.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our bounds in comparison to the existing state-of-the-art bounds. We first provide a visual of the trends of our bounds as more data is utilized for posterior predictions, as well as demonstrating the improvements that DKL can provide over standard GPs. We then demonstrate how our bounds perform in comparison to the state-of-the-art across several dynamical systems of interest. Finally, we show how our bounds can be used in safety certification, through the use of stochastic barrier functions, as described in Section 4. We note that deep kernels in general are only positive semi-definite, which prohibits the use of techniques in (Maddalena, Scharnhorst, and Jones 2021) for generating a deterministic bound. Further detail on the models considered can be seen in Appendix C.
@@
Safety Certification of Unknown Stochastic Systems We consider the Inverted Pendulum (2D) agent from the Gymnasium environment and a contractive linear 4D system for data-driven safety certification using the formulation in Section 4. We collected data for the Pendulum model under the best controller available from the OpenAI Leaderboard (OpenAI 2024) and perturb the systems with PERT distributions (a transformation of a Beta distribution with a specified mean). We construct a barrier using the Dual-Approach suggested in (Mazouz et al. 2024) and compare results when generating interval bounds on $f$ with Corollary 1 using Theorem 2, Lemma 2, and Theorem 1 with $\delta=0.05$ using DKL models.
@@ 6 Conclusion
In this paper, we derive novel error bounds (both probabilistic and deterministic) for GP regression under bounded support noise. We demonstrate that by assuming sample noise has zero mean, error bounds that are tighter than sample noise can be achieved. We show that our error bounds utilize significantly more information about the kernel than existing probabilistic bounds (i.e. each term informed by $K_{\mathcal{X},x}$); hence, they are well-suited for regression techniques based on informed kernels such as DKL that correlate $x$ with the entire dataset. We also show the improvements that our bounds can provide over the state-of-the-art in safety certification of control systems through the use of stochastic barrier functions, generating certificates with significantly larger safety probabilities. Future directions include application of these bounds in safe control and shield design in reinforcement learning.
7 Acknowledgments
This work was supported by the Air Force Research Lab (AFRL) under agreement number FA9453-22-2-0050.
Appendix A Appendix: Alternate RKHS Bounds
@@
Alternate Deterministic Error Bounds
The bounds proposed in (Maddalena, Scharnhorst, and Jones 2021) are as follows, written in a format consistent with GPR rather than Kernel Ridge Regression (recall that this work requires the kernel to be strictly positive definite). We note that this bound is larger than the bound derived in Theorem 2 in general.
Lemma 6 ((Maddalena, Scharnhorst, and Jones 2021, Theorem 1)).
Let $K_{\mathcal{X},\mathcal{X}}$ be a positive definite kernel matrix. Then with $m$ data points in $\mathcal{X}$, with observations $Y$, and $\sigma_{n}>0$, then
$|\mu_{D}(x)-f(x)|\leq P(x)\sqrt{B^{2}+\Delta-Y^{T}K_{\mathcal{X},\mathcal{X}}^{-1}Y}+\\ \tilde{\Lambda}_{x}+|Y^{T}(K_{\mathcal{X},\mathcal{X}}+\frac{1}{m\sigma_{n}}K_{\mathcal{X},\mathcal{X}}^{2})^{-1}K_{\mathcal{X},x}|$ (19)
where
$P(x)=\sqrt{\kappa(x,x)-K_{x,\mathcal{X}}K_{\mathcal{X},\mathcal{X}}^{-1}K_{\mathcal{X},x}},$
$\Delta=\max_{-\sigma_{v}\leq v_{i}\leq\sigma_{v},i=1,...,m}{\mathcal{V}^{T}K_{\mathcal{X},\mathcal{X}}^{-1}\mathcal{V}+2Y^{T}K_{\mathcal{X},\mathcal{X}}^{-1}\mathcal{V}},$
and $\tilde{\Lambda}_{x}=\sum_{i=1}^{m}\sigma_{v}|[K_{\mathcal{X},\mathcal{X}}^{-1}K_{\mathcal{X},x}]_{i}|.$
We see that $\tilde{\Lambda}_{x}$ follows a similar form to the $\Lambda_{x}$ derived in Theorem 2 (denoted as $\tilde{\Lambda}_{x}$ to make a visual comparison) however, since the inverse does not include the regularization term $\sigma_{n}^{2}I$ (i.e. $K_{\mathcal{X},\mathcal{X}}^{-1}$ versus $(K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{2}I)^{-1}$) then $\tilde{\Lambda}_{x}>\Lambda_{x}$ in general. There is also an additional additive term $|Y^{T}(K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{-2}K_{\mathcal{X},\mathcal{X}}^{2})^{-1}K_{\mathcal{X},x}|$ in the bound derived by (Maddalena, Scharnhorst, and Jones 2021), which introduces more conservatism when compared to our bound. Hence, in general their deterministic error bound will be larger than our deterministic error bound.
Work (Maddalena, Scharnhorst, and Jones 2021) states that $\Delta-Y^{T}K_{X,X}^{-1}Y$ to can be set to 0 to avoid needing to compute $\Delta$. Similarly in our setting we can set $c^{*}=0$. Note that if $\sigma_{n}=0$ then $\sigma_{D}(x)=P(x)$. Our Theorem limit $\sigma_{n}>0$ in order to guarantee the positive definiteness of $K_{X,X}+\sigma_{n}^{2}I$, however in the case that $K_{X,X}$ is strictly positive definite and invertible we can set $\sigma_{n}=0$. Then the formulation for our deterministic error bound would be written as (using the same notation as the Equation (19))
$|\mu_{D}(x)-f(x)|\leq P(x)B+\tilde{\Lambda}_{x}$ (20)
whereas Equation (19) still retains the additional additive term $|Y^{T}(K_{\mathcal{X},\mathcal{X}}+\frac{1}{m\sigma_{n}}K_{\mathcal{X},\mathcal{X}}^{2})^{-1}K_{\mathcal{X},x}|$. Note that $\sigma_{n}$ cannot be 0 in this formulation and is not typically feasible for our formulation either. However, it is clear that our bound is tighter.
Appendix B Appendix: Proofs and Extended Discussion
@@
On $\lambda_{x}$ Tends to 0 with Increased Data
Recall that $\lambda_{x}=4\sigma_{v}^{2}K_{x,\mathcal{X}}G^{2}K_{\mathcal{X},x}$ and $G=(K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{2}I)^{-1}$.
Let $\lambda_{i}$ be the $i$-th eigenvalue of $K_{\mathcal{X},\mathcal{X}}$. The eigenvalues of $G^{-1}=K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{2}I$ are then $\lambda_{i}+\sigma_{n}^{2}$. The eigen-decomposition of $G^{-1}$ is then written as $U(\Lambda+\sigma_{n}^{2}I)U^{T}$ where $U$ is an orthogonal matrix of eigenvectors and $\Lambda=\text{diag}(\lambda_{i})$. Similarly, $G=U(\Lambda+\sigma_{n}^{2}I)^{-1}U^{T}$ which has eigenvalues $1/(\lambda_{i}+\sigma_{n}^{2})$. It is known that as data is added to $K_{\mathcal{X},\mathcal{X}}$, the complexity of the matrix tends to grow which results in the eigenvalues differentiating into a cluster of larger eigenvalues and a cluster of 0 eigenvalues.
We can write $K_{\mathcal{X},x}=U\alpha$, where $\alpha$ represents the projection onto the eigenspace, which then makes the quadratic form
$K_{x,\mathcal{X}}GK_{\mathcal{X},x}=\alpha^{T}(\Lambda+\sigma_{n}^{2}I)^{-1}\alpha=\sum_{i=1}^{m}\frac{\alpha_{i}^{2}}{\lambda_{i}+\sigma_{n}^{2}}.$ (33)
We also note that it is well known that $K_{x,\mathcal{X}}GK_{\mathcal{X},x}\leq\kappa(x,x)$ for any $\sigma_{n}>0$ and any $m$ as the variance prediction of a GP cannot be negative. This directly implies that the projection of $K_{\mathcal{X},x}$ onto the eigenvectors of $G$ have minimal to zero value when $\lambda_{i}$ is close to 0 (i.e. $\alpha_{i}$ is small or zero) as this is the only way in which the sum is guaranteed to have a finite upper bound. This is due to the 0 eigenvalues of $K_{\mathcal{X},\mathcal{X}}$ being associated with eigenvectors that represent redundancy in data (or random noise in high-dimensional systems), hence they have minimal impact on the value of the inverse quadratic form. Using $G^{2}$ in the quadratic form simply results in the eigenvalues being squared in the denominator of Eqn (33), which reduces the impact of large eigenvalues as well. This results in the value of $K_{x,\mathcal{X}}G^{2}K_{\mathcal{X},x}$ tending towards 0 as more data is added.
However, while the limit $\lambda_{x}\rightarrow 0$ is theoretically feasible, it is likely computationally intractable even for reasonable choices of $\sigma_{n}$ due to the conditioning of $G$ and the $\mathcal{O}(m^{3})$ operation need to invert an $m\times m$ matrix. The term $\sigma_{n}$ can be considered as a regularization term that is designed to enable the invertability of $K_{\mathcal{X},\mathcal{X}}+\sigma_{n}^{2}I$ and also limits the upper bound of eigenvalues of $G$ to $1/(\sigma_{n}^{2})$. This enables the mathematical formulation to remain sound despite the computational intractability.
Proof of Corollary 1
Proof.
Let $\overline{\epsilon}_{X^{\prime}}(\delta)=\sup_{x\in X^{\prime}}{\epsilon}(\delta,x)$, then $\epsilon(\delta,x)\leq\overline{\epsilon}_{X^{\prime}}(\delta)$ for all $x\in X^{\prime}$ by definition of supremum. When $\mathbb{P}\Big{(}|\mu_{D}(x)-f(x)|\leq\epsilon(x,\delta)\Big{)}\geq 1-\delta$ holds, the probabilistic relation is still retained when $\epsilon(\delta,x)$ is substituted with any larger value. Hence $\overline{\epsilon}_{X^{\prime}}(\delta)$ is a valid substitution for any $x\in X^{\prime}$. ∎
Appendix C Appendix: Experiment Information